# Penalizing Length: Uncovering Systematic Bias in Quality Estimation Metrics

### v1

@@ Abstract
Quality Estimation (QE) metrics are vital in machine translation for reference-free evaluation and as a reward signal in tasks like reinforcement learning. However, the prevalence and impact of length bias in QE have been underexplored. Through a systematic study of top-performing regression-based and LLM-as-a-Judge QE metrics across 10 diverse language pairs, we reveal two critical length biases: First, QE metrics consistently over-predict errors with increasing translation length, even for high-quality, error-free texts. Second, they exhibit a preference for shorter translations when multiple candidates are available for the same source text. These inherent length biases risk unfairly penalizing longer, correct translations and can lead to sub-optimal decision-making in applications such as QE reranking and QE guided reinforcement learning. To mitigate this, we propose two strategies: (a) applying length normalization during model training, and (b) incorporating reference texts during evaluation. Both approaches were found to effectively reduce the identified length bias.
@@ 1 Introduction
Quality Estimation (QE) metric is a statistical score that predicts the quality of machine-translated text without relying on a reference translation, making it particularly useful when human-produced references are unavailable. Broadly, existing QE methods fall into two categories: regression-based models, such as MetricX (Juraska et al., 2024) and COMET (Rei et al., 2020), which learn to assess the hypothesis directly from annotated data, and LLM-as-a-Judge approaches, such as InstructScore (Xu et al., 2023) and AutoMQM (Fernandes et al., 2023), which leverage large language models to generate error ratings that approximate human judgments.
@@
Length bias refers to the systematic tendency of evaluation models to prefer or penalize outputs depending on their length rather than their true quality. Prior research has documented such effects in related evaluation settings, including LLM-as-a-Judge frameworks (Domhan & Zhu, 2025) and reward models used in reinforcement learning (Zhao et al., 2025). However, the extent to which QE metrics suffer from length bias remains largely underexplored. This gap is consequential, as length bias can distort system comparisons, over- or under-value longer translations, and ultimately misguide both model development and data curation.
@@
In this study, we design two complementary experiments to systematically investigate the presence of length bias in QE metrics. The first experiment evaluates length invariance by concatenating high-quality, error-free segments into progressively longer translations, allowing us to assess whether metrics remain stable as input length increases. The second experiment directly compares alternative hypotheses generated from the same source, enabling us to measure whether metrics exhibit a consistent preference toward shorter or longer outputs. Our analyses reveal two concerning patterns: QE metrics tend to assign spurious errors as translations grow longer, even when no actual errors are present, and they disproportionately favor shorter hypotheses when multiple candidates exist. These findings highlight a fundamental vulnerability—QE metrics risk unfairly penalizing correct but longer translations, which can in turn misguide downstream applications such as system reranking and reinforcement learning pipelines that depend on reliable QE signals.
To address the shortcomings, we propose and evaluate two mitigation strategies aimed at reducing length bias in QE metrics. The first approach introduces length normalization during model training, encouraging regression-based QE models to disentangle translation quality from sequence length and thereby improving their robustness. The second approach supplements QE evaluation with reference-based metrics to provide an external standard that counteracts length bias toward shorter outputs. Our experiments show that both strategies are effective in reducing length sensitivity and yield more reliable and equitable evaluations across translations of varying lengths.
@@ 3.2 Bias in Source Text Length
To examine bias in source text length, we construct a dataset from the WMT24++ data as follows: for each language pair, we collect the first 5 segments from each document in the data, denoted as $s_{1}$, $s_{2}$, $s_{3}$, $s_{4}$ and $s_{5}$, where each segment is approximately 1-2 sentences long. Then, starting from $s_{1}$, we gradually append the next segment $s_{i}$ to the previous segments to form passages of different lengths. Formally, for each document $d$, we construct a group of 5 passages, denoted as $p_{1}$, $p_{2}$, $p_{3}$, $p_{4}$ and $p_{5}$, where $p_{i}=s_{1}+\dots+s_{i}$.
We construct the dataset with two considerations. First, selecting segments from the same document keeps the constructed passages within a single domain and minimizes confounds from topic variation. Second, to test our hypothesis, the dataset must include source–translation pairs spanning a wide range of lengths; thus, we concatenate consecutive segments to better preserve inter-segment fluency and reduce boundary-induced disfluency that could potentially depress performance.
Since the metrics inherit different context-window limits from their base models (e.g., 512 tokens for MetricX-24 QE), we standardize the context window to the smallest across metrics, and thus discard documents whose first 5 segments exceed 500 tokens using MetricX-24’s tokenizer 111google/mt5-xl. We then evaluate each group of 5 passages using distinct metrics and analyze the relationship between source text length and translation score. To control for score variations between language pairs, we normalize scores within each group by subtracting the score of $p_{1}$ from $p2$, $p3$, $p4$, and $p5$, and report the score differences instead of raw scores. Figure 2 presents the trend of performance change with line charts over passage index across language pairs from all metrics.
@@ Results.
As shown in Figure 2, we observe a consistent trend across all language pairs: translation quality deteriorates as source text length increases. Specifically, extending the source text from 1 to 5 segments results in score drops of 1.0–2.0 points for MetricX-24 QE, 1.5–5.6 points for CometKiwi, 1.7–5.3 points for Gemini AutoMQM, and 1.8–8.5 points for Gemini LLM-as-a-Judge.
@@
To better quantify and understand this behavior, we note that the metrics are trained on MQM-annotated data using the error scheme of Freitag et al. (2021), which assigns penalties of –1 for minor errors and –5 for major errors. Consequently, when source text length increases to five segments, the metrics incorrectly predict an additional minor or major error in these error-free passages, revealing an undesired sensitivity.
@@ 4.1 Distribution of MQM Scores in WMT Data
4.2 Mitigation: Normalization at Training
To address the above length biases, we introduce length normalization, which corrects the skew in the relative frequencies of output scores. With normalization, the model is trained to predict the error density of a hypothesis translation, defined as
$D(x,h)=\frac{R(x,h)}{|h|}$
where $R(x,h)$ is the error rating of hypothesis $h$ for source $x$ and $|h|$ denotes its segment length. At inference time, we rescale the prediction back to an error rating via
$\hat{R}(x,h)=D(x,h)\cdot|h|$
@@ 4.2 Mitigation: Normalization at Training
to enable comparison with the unnormalized baseline. Figure 6 shows the gold score distributions with normalization, demonstrating greater consistency across test sets of varying lengths.
@@
To further examine the effectiveness of normalization, we trained a MetricX QE model with the above training-time normalization. Figure 6 shows the distribution of output scores using this trained model. We then replicate the experimental setup in Section 3.2 and report results in Figure 7. To highlight the effect of normalization, we compare the absolute average score difference across segment lengths (i.e., the slope of score changes). Results indicate that applying normalization reduces length bias beyond three segments, consistent with our earlier analysis of output score distributions.
4.3 Mitigation: Reference-Based Metrics
@@
When references are available, an alternative solution to mitigate evaluation-time length bias is to incorporate reference-based metrics. Figure 8 compares three classes of MT metrics: reference-based, source-based (QE), and hybrid, under the experiment setup described in Section 3.2. We observe that metrics with access to references exhibit attenuated length bias, plausibly because the reference translation provides an implicit prior on the expected length of a well-formed hypothesis. Nonetheless, these metrics continue to exhibit residual length bias, indicating that reference access alone does not fully eliminate the effect.
@@
5 Related Work
5.1 Length Effects in MT Evaluation Metrics
Early reference-based MT metrics have long been sensitive to translation length. For instance, BLEU introduces a brevity penalty (BP) to discourage overly short hypotheses that could artificially inflate n-gram precision (Papineni et al., 2002). Likewise, chrF uses an F-score over character-n-gram precision and recall, penalizing outputs that are too short (low recall) or too long (lower precision via extra content or mismatches) relative to the reference (Popović, 2015). Nonetheless, length remains a difficult dimension to handle. Papineni et al. (2002) note that neither the brevity penalty nor modified precision alone fully enforces appropriate length matching. Moreover, recent work at the document level shows that BLEU can mask quality improvements when output length increases: for example, Sun et al. (2025) find that translating entire documents produced better human-judged quality, yet BLEU scores often did not reflect this improvement, and sometimes ranked shorter, less coherent translations higher.
Beyond reference-based metrics, learned neural metrics and LLM-based evaluators also display length sensitivity. For example, Karpinska et al. (2022) reveal that even state-of-the-art metrics such as COMET can struggle with subtle perturbations, including omissions or repetitions, which may occur more frequently in unusually short or long translations. For document-level translations, Domhan & Zhu (2025) show that when using LLM-as-a-Judge translation evaluators, longer input texts tend to result in fewer detected error spans and lower system-ranking accuracy, indicating a bias related to text length rather than translation quality itself.
These findings underscore that handling translation length properly remains an important and under-addressed challenge in MT evaluation. In our work, we therefore situate ourselves within this context by investigating whether current quality-estimation (QE) methods penalize translations simply for being longer (or shorter) rather than being inherently worse.
5.2 Biases in Quality Estimation Models
Modern QE models have evolved from feature-based frameworks such as QuEst (Specia et al., 2013) to neural architectures that achieve strong results on recent benchmarks (Rei et al., 2022; 2023), with representative examples including OpenKiwi, TransQuest, and CometKiwi (Kepler et al., 2019; Ranasinghe et al., 2020; Rei et al., 2022). Despite these advances, numerous studies have noted that learned QE models often pick up spurious correlations or biases from their training data. For instance, Behnke et al. (2022) identify a partial-input bias: when trained on standard data, QE models may over-rely on features of the target translation (e.g., fluency, grammaticality) rather than verifying adequacy or fidelity to the source, leading them to assign high quality scores to fluent outputs that fail to preserve meaning.
6 Conclusion
In this work, we show that contemporary QE models lack length invariance when assessing long-form translations: they inflate error predictions as translations grow longer and exhibit a systematic bias toward shorter candidates from the same source. We trace this bias to training distributions that under-represent long, clean examples, causing models to overfit to shorter, noisier samples that dominate existing datasets. To address this issue, we propose two complementary remedies: (i) a training-time length-normalization objective that discourages sensitivity to sequence length, and (ii) an evaluation-time combination with reference-based metrics when references are available. Together, these interventions reduce length bias and move QE toward a more robust, length-agnostic assessment of translation quality. More broadly, our findings underscore the importance of accounting for input-length sensitivity in automatic evaluation systems, particularly for long-context generation tasks such as summarization and dialogue.
7 Statements
7.1 Ethics statement
We do not foresee any ethical concerns or potential risks arising from this work. All evaluation metrics and datasets are either open-sourced or accessible through public APIs, as described in Section 3.1. The large language models (LLMs) used in our experiments are also publicly available. Within our research context, the outputs generated (i.e., numerical scores from metrics and error ratings from LLMs) do not contain or propagate harmful, unsafe, or sensitive information. Accordingly, we believe that this study fully adheres to the Code of Ethics.
7.2 Reproducibility statement
We have taken several steps to ensure the reproducibility of our work. All evaluation metrics and datasets used in our experiments are open-sourced or publicly accessible. The procedures for constructing the experimental data are described in detail in Sections 3.2 and 3.4, with additional clarifications provided in the appendix. Upon acceptance, we will release the scripts necessary to replicate all experiments, including data processing, model evaluation, and result reporting. Together, these resources are intended to enable independent verification and extension of our findings.
7.3 LLM Usage Disclosure
The authors acknowledge the use of the Large Language Model Gemini 2.5 Pro for the purpose of language editing and polishing. The model was used to refine grammar, syntax, and phrasing to enhance the readability of the manuscript. The research, experimental design, and conclusions presented herein are the original work of the authors.
Appendix A Appendix
A.1 Proportion of Scores with Decreasing Trends in Data Filtering

### v2

@@ Abstract
Quality Estimation (QE) metrics are vital in machine translation for reference-free evaluation and as a reward signal in tasks like reinforcement learning. However, the prevalence and impact of length bias in QE have been underexplored. Through a systematic study of top-performing regression-based and LLM-as-a-Judge QE metrics across 10 diverse language pairs, we reveal two critical length biases: First, QE metrics consistently over-predict errors with increasing translation length, even for high-quality, error-free texts. Second, they exhibit a preference for shorter translations when multiple candidates are available for the same source text. These inherent length biases risk unfairly penalizing longer, correct translations and can lead to sub-optimal decision-making in applications such as QE reranking and QE guided reinforcement learning. We further investigate the root cause, presenting evidence that QE models conflate quality with sequence length due to skewed supervision distributions. As a diagnostic intervention, we apply length normalization during training. We show that this simple intervention is sufficient to decouple error probability from length, effectively counteracting the data skew and yielding more reliable QE signals.
Machine Translation (MT), Quality Estimation, LLM-as-a-Judge
@@ 1 Introduction
Quality Estimation (QE) predicts the quality of machine-translated text without relying on reference translations, serving as a vital signal for applications where ground truth is unavailable. Broadly, existing QE methods fall into two categories: regression-based models, such as MetricX (Juraska et al., 2024) and COMET (Rei et al., 2020), which learn to assess the hypothesis directly from annotated data, and LLM-as-a-Judge approaches, such as InstructScore (Xu et al., 2023) and AutoMQM (Fernandes et al., 2023), which leverage large language models to generate error ratings that approximate human judgments.
@@
Length bias refers to the systematic tendency of evaluation models to prefer or penalize outputs depending on their length rather than their true quality. Prior research has documented such effects in related evaluation settings, including LLM-as-a-Judge frameworks (Domhan and Zhu, 2025) and reward models used in reinforcement learning (Zhao et al., 2025). However, the extent to which QE metrics suffer from length bias remains largely underexplored. This gap is consequential, as length bias can distort system comparisons, over- or under-value longer translations, and ultimately misguide both model development and data curation.
@@
In this study, we design two complementary experiments to systematically investigate the presence of length bias in QE metrics. The first experiment evaluates length invariance by concatenating high-quality, error-free segments into progressively longer translations, allowing us to assess whether metrics remain stable as input length increases. The second experiment directly compares alternative hypotheses generated from the same source, enabling us to measure whether metrics exhibit a consistent preference toward shorter or longer outputs. Our analyses reveal two concerning patterns: QE metrics tend to assign spurious errors as translations grow longer, even when no actual errors are present, and they disproportionately favor shorter hypotheses when multiple candidates exist. These findings highlight a fundamental vulnerability: QE metrics exhibit a systematic length penalty, assigning spurious errors to longer segments and disproportionately favoring shorter hypotheses in reranking scenarios.
To address these shortcomings, we analyze the training data of a representative state-of-the-art QE metric, MetricX-24, and identify an implicit label imbalance. Specifically, supervision signals are unevenly distributed across sequence lengths, inducing a spurious association whereby longer translations are more likely to contain errors and thus receive worse labels. Motivated by this diagnosis, we apply length normalization during training to counteract this length-dependent skew. Experiments show that length-normalized training substantially reduces length sensitivity and yields more reliable QE signals across translations of varying lengths.
2 Related Work
2.1 Length Effects in MT Evaluation Metrics
Early reference-based MT metrics have long been sensitive to translation length. For instance, BLEU introduces a brevity penalty (BP) to discourage overly short hypotheses that could artificially inflate n-gram precision (Papineni et al., 2002). Likewise, chrF uses an F-score over character-n-gram precision and recall, penalizing outputs that are too short (low recall) or too long (lower precision via extra content or mismatches) relative to the reference (Popović, 2015). Nonetheless, length remains a difficult dimension to handle. (Papineni et al., 2002) note that neither the brevity penalty nor modified precision alone fully enforces appropriate length matching. Moreover, recent work at the document level shows that BLEU can mask quality improvements when output length increases: for example, (Sun et al., 2025) find that translating entire documents produced better human-judged quality, yet BLEU scores often did not reflect this improvement, and sometimes ranked shorter, less coherent translations higher.
Beyond reference-based metrics, learned neural metrics and LLM-based evaluators also display length sensitivity. For example, Karpinska et al. (2022) reveal that even state-of-the-art metrics such as COMET can struggle with subtle perturbations, including omissions or repetitions, which may occur more frequently in unusually short or long translations. For document-level translations, (Domhan and Zhu, 2025) show that when using LLM-as-a-Judge evaluators, longer input texts tend to result in fewer detected error spans and lower system-ranking accuracy, indicating a bias related to text length rather than translation quality.
These findings underscore that handling translation length properly remains an important and under-addressed challenge in MT evaluation. In our work, we therefore situate ourselves within this context by investigating whether current QE methods penalize translations simply for being longer (or shorter) rather than being inherently worse.
2.2 Biases in Quality Estimation Models
Modern QE models have evolved from feature-based frameworks such as QuEst (Specia et al., 2013) to neural architectures that achieve strong results on recent benchmarks (Rei et al., 2022, 2023), with representative examples including OpenKiwi, TransQuest, and CometKiwi (Kepler et al., 2019; Ranasinghe et al., 2020; Rei et al., 2022). Despite these advances, numerous studies have noted that learned QE models often pick up spurious correlations or biases from their training data. For instance, (Behnke et al., 2022) identify a partial-input bias: when trained on standard data, QE models may over-rely on features of the target translation (e.g., grammaticality) rather than verifying adequacy or fidelity to the source, leading them to assign high quality scores to fluent outputs that fail to preserve meaning.
@@ 4.2 Bias in Source Text Length
To examine bias in source text length, we construct a dataset from the WMT24++ data as follows: for each language pair, we collect the first 5 segments from each document in the data, denoted as $s_{1}$, $s_{2}$, $s_{3}$, $s_{4}$ and $s_{5}$, where each segment is approximately 1-2 sentences long. Then, we reconstruct the original document flow by gradually appending the subsequent contiguous segment $s_{i}$ to the previous sequence. Formally, for each document $d$, we construct a group of 5 passages, denoted as $p_{1}$, $p_{2}$, $p_{3}$, $p_{4}$ and $p_{5}$, where $p_{i}=s_{1}+\dots+s_{i}$. This ensures that the resulting passages are not arbitrary concatenations, but represent the original, coherent discourse intended by the source texts.
We construct the dataset as such with two considerations. First, selecting segments from the same document keeps the constructed passages within a single domain and minimizes confounds from topic variation. Second, to test our hypothesis, the dataset must include source–translation pairs spanning a wide range of lengths; thus, we concatenate consecutive segments to preserve inter-segment fluency better and reduce boundary-induced disfluency that could potentially depress performance.
Since the metrics inherit different context-window limits from their base models (e.g., 512 tokens for MetricX-24 QE), we standardize the context window to the smallest across metrics, and thus discard documents whose first 5 segments exceed 500 tokens using MetricX-24’s tokenizer  [^1] https://huggingface.co/google/mt5-xl. We then evaluate each group of 5 passages using distinct metrics and analyze the relationship between source text length and translation score. To control for score variations between language pairs, we normalize scores within each group by subtracting the score of $p_{1}$ from $p2$, $p3$, $p4$, and $p5$, and report the score differences instead of raw scores. Figure 2 presents the trend of performance change with line charts over passage index across language pairs from all metrics.
@@ Results.
As shown in Figure 2, we observe a consistent trend across all language pairs: translation quality deteriorates as source text length increases. Specifically, extending the source text from 1 to 5 segments results in score drops of 4.0–8.0 points for MetricX-24 QE, 1.5–5.6 points for CometKiwi, 1.7–5.3 points for Gemini AutoMQM, and 1.8–8.5 points for Gemini LLM-as-a-Judge.
@@
To better quantify this behavior, we note that the metrics utilize the MQM error scheme (Freitag et al., 2021), which assigns penalties of –1 (minor) and –5 (major) on a 25-point scale; normalized to our 100-point scale, these correspond to penalties of –4 and –20, respectively. While a marginal decrease in quality scores might be expected due to the cumulative probability of minor imperfections in longer sequences ($P^{N}_{error-free}$), the magnitude of the observed drop, particularly in LLM-as-a-Judge metrics (up to 8.5 points), far exceeds the theoretical decay of joint error probability. This suggests the penalty is driven by a systematic bias against length rather than a valid accumulation of errors.
@@ 5.1 Distribution of MQM Scores in WMT Data
5.2 Length Normalization at Training
To validate our hypothesis that the bias stems from label density skew rather than architectural limitations, we apply a diagnostic intervention: simple length normalization. We posit that if a linear correction factor ($1/|h|$) is sufficient to remove the bias, the root cause is confirmed to be overfitting to the training data’s label distribution.
@@ 5.2 Length Normalization at Training
With normalization, the model is trained to predict the error density of a hypothesis translation, defined as $D(x,h)=R(x,h)/|h|$ where $R(x,h)$ is the error rating of hypothesis $h$ for source $x$ and $|h|$ denotes its segment length. At inference time, we rescale the prediction back to an error rating via $\hat{R}(x,h)=D(x,h)\cdot|h|$ to enable comparison with the unnormalized baseline. Figure 6 shows the gold score distributions with normalization, demonstrating greater consistency across test sets of varying lengths.
To further examine the effectiveness of normalization, we trained a MetricX QE model with the above training-time normalization. Figure 6 (Bottom-Left) demonstrates the impact of this intervention. Crucially, the normalized model’s score distribution mirrors the ground-truth human distribution (Figure 6, Bottom-Right): the curves for different lengths collapse onto a single, invariant density function. This confirms that length normalization does not merely dampen scores, but successfully recalibrates the model to match the fundamental statistical properties of human quality judgments—a property the unnormalized baseline fails to capture. By restoring this distributional alignment, the model learns to decouple error probability from sequence length.
@@
Additionally, we replicate the experimental setup in Section 4.2 and report results in Figure 7. To highlight the effect of normalization, we compare the absolute average score difference across segment lengths (i.e., the slope of score changes). Results indicate that applying normalization reduces length bias beyond three segments, consistent with our earlier analysis of output score distributions.
6 Conclusion
In this work, we show that contemporary reference-free QE models are not length-invariant when assessing long-form translations: they inflate predicted error as translations grow longer and systematically favor shorter candidates for the same source. We trace this phenomenon to a training pipeline artifact: the under-representation of long, error-free examples in standard datasets. By demonstrating that a simple length-normalization step restores alignment with human ground-truth distributions, we provide a diagnostic proof that the bias is distributional rather than architectural. While standard aggregate benchmarks often miss these tail-end failures, our findings highlight the necessity of explicitly auditing length-invariance to ensure reliable evaluation in long-context applications.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Translation evaluation. Beyond technical improvements, our findings have potential implications for linguistic fairness and model safety. We demonstrate that current QE metrics systematically penalize length, which may result in the unfair devaluation of translations in languages that are naturally more verbose (e.g., German, Russian) compared to concise languages (e.g., English, Chinese). Furthermore, as QE metrics are increasingly employed as reward signals in Reinforcement Learning (RLHF), unmitigated length bias poses a risk of driving generation models toward excessive brevity, potentially leading to information loss or over-simplification of complex source concepts. By diagnosing and mitigating this bias, our work aims to ensure that automated evaluation remains equitable across diverse linguistic typologies and promotes reliability in automated training pipelines.
Appendix A Appendix
A.1 Length Bias in Reference-Based Metrics
@@
When references are available, we compare reference-based, source-based (QE), and hybrid MT metrics under the experiment setup described in Section 4.2. Figure 8 shows that metrics with access to references exhibit attenuated length bias, plausibly because the reference translation provides an implicit prior on the expected length of a well-formed hypothesis. Nonetheless, these metrics continue to exhibit residual length bias, indicating that reference access alone does not fully eliminate the effect.
@@
A.2 Length Bias with respect to Model Size
Some of the QE models we study (e.g., CometKiwi and LLM-as-a-Judge) are only available in a single model size. In contrast, MetricX-24 provides multiple variants (Large, XL, and XXL). In this subsection, we examine how length bias varies with model capacity by repeating our main diagnostics using these MetricX-24 variants.
Experimental setup.
Unless otherwise stated, we use the bfloat16 variants of the Large and XL models for consistency with the publicly available bfloat16 XXL checkpoint. [^2] https://huggingface.co/google/metricx-24-hybrid-xxl-v2p6-bfloat16
Experiment 1: Bias in source/translation length.
We repeat the concatenation experiment from Section 3.2 using all three MetricX-24 variants. Table 1 reports the average predicted error as we concatenate five error-free segments. We observe a clear relationship between model size and the severity of the length-related penalty: smaller models exhibit stronger bias.
Model 1 2 3 4 5
Large 0.00 1.41 2.49 3.38 4.16
XL 0.00 1.02 1.61 2.24 2.61
XXL 0.00 0.55 0.80 1.26 1.47
Table 1: Model-size analysis on concatenating error-free segments (passage index 1155). Smaller MetricX-24 variants exhibit larger increases in predicted error as length grows.
Experiment 2: Bias in translation length preferences.
We repeat the reranking-style length preference experiment from Section 3.4. In an unbiased setting, a model should show roughly 50% preference for the shorter translation when two candidates are of comparable quality but differ in length. Table 2 reports the preference rates for the shorter translation at different length gaps. Consistent with Experiment 1, the smaller model exhibits a more pronounced length preference, while the XL and XXL variants behave similarly.
Model 2.5% 5% 7.5% 10% 12.5% 15%
Large 57.2 57.3 57.3 57.7 59.6 62.5
XL 53.5 53.6 53.4 54.4 55.9 56.5
XXL 54.3 54.0 53.4 54.0 56.8 57.8
Table 2: Model-size analysis on length preference: percentage preference for the shorter translation under different length differences. Values closer to 50% indicate less length bias.
Taken together, these results suggest that increasing model capacity can reduce length bias, but substantial bias remains even for the largest publicly available MetricX-24 variant, motivating model-agnostic mitigation strategies such as length-normalized training.
A.3 Proportion of Scores with Decreasing Trends in Data Filtering